MIDC News
Year of 2012
December 18, 2012
- 08:56 MST: Due to data logger clock corruption, SMUD Anatolia data from Dec 17, 2012 has been removed from MIDC until further notice. -Afshin Andreas
- 14:26 MST: The Natural Energy Laboratory of Hawaii Authority (NELHA) measurement station has been added to MIDC. Data is available since November 1, 2012.
-Afshin Andreas - 10:31 MST: Due to data logger clock corruption, SMUD Anatolia data from Dec 14 through 16, 2012 has been removed from MIDC until further notice. -Afshin Andreas
- 10:03 MST: SRRL BMS historical preset equations used from 1985 to 2003 were updated to use the correct input variables, which were never updated for a revised preset configuration that was made several years after. -Afshin Andreas
- 10:46 MST: The preset equations which are available for some measurement stations (which are calculated on the fly from raw data), are now defined for the user. Simply click the word "Preset" next to the radio button that appears on the download page for an applicable station. -Afshin Andreas
- 11:19 MST: At 9:40 MST today the SRRL BMS data stream was upgraded to version 12.0. All historical data since June 1, 2012 was converted to this format. The following changes occured:
- Add raw mV signal output for all pyranometers which have a net infrared thermal correction applied
- Removal of two Global PSP readings with no zenith angle correction
- Removal of Diffuse PSP with no net infrared correction
- Add Direct Quantum LI-190 sensor
- Rename CNR1 calculated fields to K&Z and use Downwelling CG4 and CM22 instruments instead of Downlwelling CG3 and CM3 for Albedo, Atmos Net Infrared, and Net Radiation.
- Restructure raw data field order to support future instrument re-location and new radiometer tower.
- 08:50 MST: No new SRRL PWV processed data has appeared on the UCAR website since November 19, 2012 at 13:45 MST. However, Trimble at SRRL is still collecting raw data, problem is unknown at this time.
-Afshin Andreas - 08:49 MST: Communication with the NREL VTIF RSR was lost on Nov 20 @ 10:56 MST. Problem is unknown at this time. -Afshin Andreas
- 08:48 MST: There has been a power outage at SolarTAC since 11/20/2012 at 16:46 MST. Therefore solar tracker is not operating (direct and diffuse data is bogus). -Afshin Andreas
- 09:42 MST: The SRRL BMS Upwelling PIR was troubleshot and disconnected from service yesterday, problem most likely due to water infiltration into j-box and signal cable to instrument. Data is bogus since Nov 9 at 16:30 MST, for the Upwelling PIR, Net Infrared PIR (atmos), and Net Radiation (Eppley). -Afshin Andreas
- 17:11 MST: The SRRL BMS Met/PV Tower CR10X logger was upgraded to a table based OS at 16:16 MST today. The same was done for the Radiometer Tower CR10X logger at 16:42 MST today. -Afshin Andreas
- 08:01 MST: The SRRL BMS Instrument Deck data logger was upgraded from a CR23X to a CR3000 on November 9 at 16:30 MST. -Afshin Andreas
- 12:03 MST: The OTF Global 40-South MS-700 Spectroradiometer was removed from service on Oct 16 @ 13:05 MST, calibrated against NIST F655, and returned to service today @ 8:14 MST. -Afshin Andreas
- 08:52 MST: The SRRL PGS-100 spectroradiometer was calibrated on September 28, 2012 outdoors against the NIST traceable LI-1800 PRS-102. The new calibration file will be applied to data starting at 8:35 MST today. -Afshin Andreas
- 15:26 MST: Solar and Moon Position Algorithm (SAMPA) Calculator has been developed for MIDC. -Afshin Andreas
- 15:50 MST: The MIDC primary data acquisition PC was upgraded to Dell Precision 390 Core 2 Quad with SSD at 13:51 MST today. -Afshin Andreas
- 08:33 MST: The SRRL TSI-880 shading band was repaired at 08:32 MST today. The SEI bus to RS-232 converter for the mirror encoder was replaced (model AD2-B). -Afshin Andreas
- 15:28 MST: The MIDC primary data acquisition was restored at 13:31 MST today. -Afshin Andreas
- 06:33 MST: The MIDC primary data acquisition system (which collects data from all "real time stations", such as SRRL BMS, NWTC M2, VTIF, SolarTAC, SSRP, NREL Spectral Data) failed at 1:00 am this morning due to a corrupt system file. Therefore data collection for those stations are currently down until this is resolved.
-Afshin Andreas
- 09:21 MST: The temperature, wind speed, wind direction, and dew point sensors at NWTC M2 were replaced with calibrated instruments on August 15, 2012. The data logger program was updated today @ 8:51 MST with new calibration constants for the wind speed and temperature sensors. -Afshin Andreas
- 14:44 MST: The following SRRL BMS Radiometer Tower instruments were swapped out with instruments calibrated in BORCAL 2012-01 between 13:52 and 14:17 MST today: Global 90-North/East/South/West PSP & LI-200, Global 40-South PSP & LI-200, Upwelling PSP & LI-200, and the Global Normal CM21. Calibration factors were updated in the logger @ 14:09 MST today. -Afshin Andreas
- 15:51 MST: The SRRL Global 40-South LI-1800 Spectroradiometer was removed from service today @ 13:57 MST, calibrated against NIST F655, and returned to service @ 15:47 MST. -Afshin Andreas
- 15:26 MST: The following SRRL BMS pyrgeometers were swapped out today with calibrated instruments between 14:50 and 14:59 MST: Downwelling PIR, Upwelling PIR, and Downwelling CG4. New calibration coefficients were updated in the logger at 14:58. -Afshin Andreas
- 07:55 MST: The shading band tracking system on the SRRL TSI-880 total sky imager failed on 8-2-2012 at approx 17:00 MST, however images and data are still being collected. Problem is unknown at this time. -Afshin Andreas
- 10:26 MST: All SRRL BMS UV instruments were calibrated on 06-13-2012, this includes: UV-S-B-T, CUVB1, MS210W, UVB-1, 501A, TUVR, CUV4 (both), UV-S-A-T, CUVA1, CUVA2, CUVB2. Logger was updated with new cal factors today @ 10:19 MST. The UV-S-B-T appears to have a problem since its calibration factor changed by 15% and it has already changed after only 1 month. -Afshin Andreas
- 08:05 MST: The SRRL TSR-1's and the MFR-7 were permanently shutdown yesterday. The evaluation period has ended and they will be retured to YES. -Afshin Andreas
- 11:40 MST: The power supply in the SRRL Total Sky Imager was replaced today. Cloud cover and images are available as of 11:12 MST. -Afshin Andreas
- 09:57 MST: The SRRL Global LI-1800 spectroradiometer is working as of 9:40 MST today. -Afshin Andreas
- 10:46 MST: The SRRL Global LI-1800 Spectroradiometer is down since July 6 at 15:20 MST. Likely due to thunderstorm activity, instrument may have failed and will be down until further notice. -Afshin Andreas
- 07:53 MST: The value of ΔT has been incorrectly computed with the equation: ΔT = 32.184 + (TAI-UTC) + ΔUT1. The value was corrected today at 7:49 MST using the correct equation of ΔT = 32.184 + (TAI-UTC) - ΔUT1. -Afshin Andreas
- 12:01 MST: There was a power glitch at NREL at 1:31 MST, as a result the SRRL TSI-880 total sky imager is dead. Therefore images and cloud cover data is not available until further notice. -Afshin Andreas
- 09:17 MST: The SRRL BMS Global SE-POA LI-200 was swapped out with a sensor calibrated in BORCAL 2012-01 at 8:52 MST today. Cal factor was updated in the logger @ 8:59 MST today. -Afshin Andreas
- 09:02 MST: The shadow band correction factor (SBCF) was removed from the SRRL BMS shadow band PSP on 6/13/2012 at 12:34 MST. The instrument appears to be over corrected now that the thermal correction has been applied. Note: For research purposes the data w/SBCF is being stored in "Research Field 0". -Afshin Andreas
- 12:39 MST: SolarTAC has been without power since June 6 at 13:55 MST. Since the solar tracker is down, direct normal and diffuse measurements are bad. -Afshin Andreas
- 12:28 MST: As of today all SRRL BMS horizontal non-black/white thermopile pyranometers (except the CM3) are corrected using the effective net infrared irradiance measured by the downwelling PIR or CG4. Update took place at 8:50 MST for these instruments: Global CM22 and CM6b; and at 10:28 MST for these instruments: Global RG780 PSP, Zebra PSP, Shadowband PSP, Diffuse CM22. Note: The Global PSP (vent and non-vent) and the Diffuse PSP already had the correction applied (any non-corrected values are no longer available as of 8:50 MST). -Afshin Andreas
- 12:22 MST: The following SRRL BMS deck instruments were swapped out with instruments that were calibrated in BORCAL 2012-01 between approx 10:00 to 10:28 MST today: Global RG780 PSP, Direct RG780 NIP, Zebra PSP, Shadowband PSP, Diffuse CM22, Diffuse PSP, and Diffuse 8-48. Cal factors were updated in the logger @ 10:28 MST today. -Afshin Andreas
- 12:18 MST: The following SRRL BMS instruments were calibrated in BORCAL 2012-01: Global Horizontals: CM22, CM6b, PSP (ventilated), PSP, TSP-700, TSP-1, SP-Lite, LI-200, SP-110, SPN1; Direct Normals: CH1, NIP #2, NIP #1, LI-201; Diffuse SPN1. Logger program was updated at 08:50 MST today with new cal factors. Responsivity Coefficients for Global PSPs were also updated at 08:54 MST today. -Afshin Andreas
- 17:17 MST: There is planned power outage for SRRL on Saturady May 26 from 7 AM until 4 PM. Data updates for all MIDC stations could be down during this time. Also several SRRL BMS instruments will also be without power during this time. -Afshin Andreas
- 13:31 MST: SolarTAC CMP22, CH1, and 8-48 were replaced with calibrated instruments today. Calibration factors were updated at 13:25 MST today. -Afshin Andreas
- 07:42 MST: Communication with Research Park was lost on May 9, 2012 at 16:57 MST. Problem is unknown at this time.
-Afshin Andreas
- 15:21 MST: An unauthorized data logger program change was made to UTPA SRL as a result CR1000 temperature is missing, irradiance values have lost their precision, and there are extra unlabeled columns at the end of the raw data stream. -Afshin Andreas
- 15:03 MST: The NREL Vehicle Testing and Integration Facility (VTIF) Rotating Shadownband Radiometer (RSR) has been added to MIDC. Data is updated every 60 seconds and is available since April 1, 2012. -Afshin Andreas
- 10:57 MST: The NWTC M2 PSP was removed from service at 10:36 MST and replaced with a calibrated instrument at 10:44 MST today. New calibration factor for S/N 20517F3 was entered into the logger at 10:53 MST today. -Afshin Andreas
- 07:50 MST: Data collection to ORNL resumed at 7:33 MST today. -Afshin Andreas
- 09:31 MST: Data collection to the ORNL RSR has been down since March 20. Restoration of communication is unknown at this time. -Afshin Andreas
- 09:25 MST: The SRRL BMS CR23x data logger was calibrated at 8:55 to 9:15 MST today. The following instruments may read bogus during that time: Diffuse Shadowband PSP, Direct LI-201, Global SP-Lite, Global SP-110, Global TSP-1. -Afshin Andreas
- 08:49 MST: Data collection for LMU resumed as of 8:37 MST today. -Afshin Andreas
- 15:48 MST: Data collection for Research Park resumed as of 15:26 MST today. -Afshin Andreas
- 15:55 MST: SRRL data lab temp/rh probe was returned to service at 15:52 MST today. -Afshin Andreas
- 10:49 MST: The SRRL data lab temperature/rh probe was removed from service at 10:03 MST today for calibration. -Afshin Andreas
- 08:13 MST: As of 08:12 MST today data collection to all stations (except Research Park and LMU) has been restored. -Afshin Andreas
- 08:14 MST: Data collection for SS2 resumed as of 4:09 MST today. -Afshin Andreas
- 09:22 MST: Data collection for the following stations has resumed as 9:20 MST today: LRSS, SMUD Anatolia, USEP Milford, USEP Cedar City, and Tri-State Escalante. -Afshin Andreas
- 08:54 MST: Data collection to SolarTAC resumed at 08:54 MST today. -Afshin Andreas
- 07:58 MST: Data collection from most non-NREL stations are currently down due to a NREL cyber security change that occurred on March 9 @ 12:03 MST, which include (but not limited to): BSC, HSU SoRMS, LRSS, SMUD Anatolia, SolarTAC, LMU, USEP Milford & Cedar City, Research Park, SunSpot2, Tri-State Escalante, UA OASIS, and UTPA SRL. Restoration of communication to these stations are unknown at this time. -Afshin Andreas
- 09:11 MST: October 2011 through February 2012 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 08:35 MST: October 2011 through February 2012Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 11:43 MST: Langley plots have been generated for all SRRL photometers and Vo values were updated today @ 11:31 MST. -Afshin Andreas
- 14:08 MST: The SRRL BMS POM-01 Photometer was returned to service at 14:00 MST. The factory upgraded the performance of the 315nm measurement. -Afshin Andreas
- 10:48 MST: The URL for the SRRL BMS and NWTC M2 PDA text based display have been greatly simplified, so that bookmarks do not become obsolete when the instrument configuration changes for those sites. Existing bookmarks should be updated to prevent future problems. -Afshin Andreas
- 14:54 MST: The SRRL PGS-100 spectroradiometer was calibrated on February 27, 2012 outdoors against the NIST traceable LI-1800 PRS-158. The new calibration file will be applied to data starting at 14:55 MST today. -Afshin Andreas
- 16:39 MST: The SRRL Global Horizontal LI-1800 spectroradiometer was moved to it's final location on the instrument deck today @ 12:30 MST. -Afshin Andreas
- 15:58 MST: The SRRL LI-1800 Global Spectroradiometer was moved from the 40-South Tilt position to the horizontal position at 15:55 MST. The instrument is temporarily located on the SRRL south patio but will be relocated to the SRRL instrument deck within the next week. -Afshin Andreas
- 08:59 MST: We are currently experiencing communications problems with our sites using Verizon cell phone modems. All sites using Verizon are currently down and data has not updated since February 1, 2012 14:53 MST. -Afshin Andreas
- 14:13 MST: The SRRL BMS Temp/RH HMP-45C probe (PV) was swapped out with a calibrated sensor today between 13:39 to 13:44 MST -Afshin Andreas
- 15:10 MST: The wireless ethernet antenna for the NTWC M2 tower data was damaged by high wind at 12:56 MST today. NWTC M2 data will not update until it has been repaired. -Afshin Andreas
- 14:46 MST: The SRRL Global 40-South LI-1800 Spectroradiometer was removed from service today @ 12:57 MST, calibrated against NIST F655, and returned to service @ 14:37 MST -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]